OpenStudy (anonymous):

what number must be added to the polynomial to complete this square? x^2-22x

OpenStudy (anonymous):

im a bit slow:/

OpenStudy (anonymous):

ok here we go.

OpenStudy (anonymous):

half of -22 is...

OpenStudy (anonymous):

slow

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

and \[(-11)^2\] is

OpenStudy (anonymous):

121

OpenStudy (anonymous):

you win!

OpenStudy (anonymous):

oh, i didn't know that rule still applied to polynomials. thank you:)

OpenStudy (anonymous):

? yes it does. btw when you add 121 you get \[x^2-22x+121=(x-11)^2\]

OpenStudy (anonymous):

Yeah, as long as x^2 term has coefficient of 1, this rule works For Example, it won't work for 2x^2+4x

OpenStudy (anonymous):

what if you would solve for x in a binomial equation?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

solving for x in a polynomial i meant

OpenStudy (anonymous):

Say that you are trying to solve for x x^2−22x+121=0 So you factor (x-11)^2=0 that means x=11

OpenStudy (anonymous):

so say x^2-7x= - 24/4 run me through how to solve it step by step.. sorry for the trouble

OpenStudy (anonymous):

please..

OpenStudy (anonymous):

are you still trying to do this problem?